Johns test scores are 95,90,83. What must he score on his next test to have an average test score of 90 on the four test?
Answer by shweta(56)

John's test scores are 95,90,83 ,x
Average test score in four tests= 90 .
We need to find the score in the fourth test, x
Average, 90= (95+ 90+ 83+x)/4
On cross- multiplication
95+90+83+x= 90*4
268+ x= 360
x= 360-268
x= 92, the score in the fourth test